OpenStudy (anonymous):

-2 ( q+8) + 2( 4q-3) = 2( 4q+3)

OpenStudy (anonymous):

Can you multiply these out? :o)

OpenStudy (anonymous):

-2 - 16 + 8q - 6 = 8q + 6

OpenStudy (anonymous):

Great! So now combine like terms: -2q-16+8q-6=8q+6 6q-22=8q+6 Do you see how I got that?

OpenStudy (anonymous):

wait how did u get that

OpenStudy (anonymous):

Well when adding like terms, you're going to add the variables and then the regular numbers without variables. So: On the left side of the equal sign: -2q-16+8q-6 ------------- -2q+8q=6q -16+(-6)= -22 Therefore: 6q-22 So, 6q-22=8q+2 The right side had no like terms to combine. Do you understand now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now, we are going to get q on the same side of the equation, so: 6q-22=8q+2 -6q -6q ^this cancels the 6q. 6q-6q=0. -22=2q+2 Do you understand that? :o)

OpenStudy (anonymous):

So, now, we're going to get q alone. So, we need to get rid of the +2 and the *2. So: -22=2q+2 -2 -2 ^This cancels. 2-2=0 -24=2q Now, 2q is multiplying, so the opposite will be to be divide. So: -24=2q /2 /2 ^This will cancel. 2/2=1. And 1q is the same as just q. Can you figure out the value of q now? :o)

OpenStudy (anonymous):

how did u get plus 2 when it was plus 6?

OpenStudy (anonymous):

12! :)

OpenStudy (anonymous):

Oops: 6q-22=8q+6 -6q -6q -22=2q+6 -6 -6 -28=2q /2 /2 -14=q